Jack will need R20000 to buy his brother's car in two year's time. He wants to start saving part of his weekly salary into an account that returns 8,5% interest per year, compounded weekly calculate the minimum weekly payment that he needs to make into this investment account to have enough in two year's time
Answer by Theo(13342)

i used the ti-ba-ii financial calculator to solve this.
inputs are:
present value (pv) = 0
future value (fv) = 20,000
number of time periods (n) = 2 years * 52 weeks per year = 104 weeks.
interest rate per time period (i/y) = 8.5% per year divided by 52 = .163461538% per week.
payments are made at the end of each time period (2d pmt) = at the end of each week.
i clicked on payment per time period (pmt) to get 176.58 rounded to the nearest penny.
the calculator says that it is 176.5812108 rounded to the number of digits available to be displayed.
the total payments = 176.5812108 * 104 = 18364.44592.
the total interest = 20,000 minus that = 1635.554081.
your solution is that the minimum deposited at the end of each week is equal to 176.5812108 which can be rounded to 176.58 if rounding to the nearest penny, or rounded to 176.59 if rounding to the nearest penny that will put at least 20,000 into the account by the end of the investment period.
the difference will be very small either way (within a few pennies at most).
you can duplicate the results by using the financial calculator at https://arachnoid.com/finance/index.html
give it a try.
if you can't do it, let me know and i'll guide you through it.
theo